OpenStudy (anonymous):

prove that : cot(90-theta)/tan theta + cosec(90-theta) x sin theta/tan(90-theta) = sec sq theta

OpenStudy (sidsiddhartha):

do u know basic degree transformation like \[\sin(90-\theta)=\cos(\theta)\] \[\tan(90-\theta)=\cot \theta\] etc

OpenStudy (anonymous):

have you tried solving it neha?

OpenStudy (goformit100):

Warm Welcome To Open Study ......... Sir/ Ma'am @nehalashafi

OpenStudy (anonymous):

what is 'sec sq theta' ?

OpenStudy (anonymous):